# Semma: A Semantic Aware Knowledge Graph Foundation Model

### v1

@@ Abstract
Knowledge Graph Foundation Models (KGFMs) have shown promise in enabling zero-shot reasoning over unseen graphs by learning transferable patterns. However, most existing KGFMs rely solely on graph structure, overlooking the rich semantic signals encoded in textual attributes. We introduce Semma, a dual-module KGFM that systematically integrates transferable textual semantics alongside structure. Semma leverages Large Language Models (LLMs) to enrich relation identifiers, generating semantic embeddings that subsequently form a textual relation graph, which is fused with the structural component. Across 54 diverse KGs, Semma outperforms purely structural baselines like Ultra in fully inductive link prediction. Crucially, we show that in more challenging generalization settings, where the test-time relation vocabulary is entirely unseen, structural methods collapse while Semma is $2$x more effective. Our findings demonstrate that textual semantics are critical for generalization in settings where structure alone fails, highlighting the need for foundation models that unify structural and linguistic signals in knowledge reasoning.
@@ 1 Introduction
In analogy to Large Language Models (LLMs) that learn complex correlations between tokens from a large corpus and apply them in previously unseen contexts Zhao et al. (2025), related work has started to investigate Knowledge Graph Foundation Models (KGFMs) to learn complex reasoning capabilities from training knowledge graphs and apply them to previously unseen test knowledge graphs Galkin et al. (2024); Zhang et al. (2025b); Huang et al. (2025).
@@
Generalizing within-knowledge graph embedding methods to transfer embeddings from one or multiple knowledge graphs with millions of nodes to an unseen knowledge graph with largely distinct nodes can be hindered by a low signal-to-noise ratio Zhao et al. (2021). However, we observe that knowledge graphs tend to have a huge number of nodes, but a much smaller number of highly informative relations. E.g., Wikidata Vrandecic and Krötzsch (2014) contains over 110 million nodes but only 12,681 relations as of May 19, 2025.
@@
As illustrated in Fig. 1, our approach constructs: (i) a structural relation graph, which represents patterns that generalize from graph structures of the training graphs, similar to Ultra (Galkin et al., 2024), and (ii) a textual relation graph, which represents patterns that are found by prompting an LLM to generate semantically rich descriptions for each relation identifier and/or attributed labels. Together, the two relation graphs allow for predicting high-quality links that would not be found by existing KGFM methods. In summary, our contributions are:
@@ 2 Preliminaries
Given the observed graph $\mathcal{G}_{\textsc{o}}$, link prediction aims to infer missing true triples $\mathcal{F}_{\textsc{t}}\setminus\mathcal{F}_{\textsc{o}}$ (called queries) using triples with either the head or tail entity masked, denoted as $(h,r,?)$ and (?$,r,t)$. We assume that for every relation, denoted by $r$, there is an inverse relation, denoted by $r^{-1}$. Let $\mathcal{G}_{\textsc{train}}=(\mathcal{E}_{\textsc{train}},\mathcal{R}_{\textsc{train}},\mathcal{F}_{\textsc{train}})$ correspond to the training graph and $\mathcal{G}_{\textsc{test}}=(\mathcal{E}_{\textsc{test}},\mathcal{R}_{\textsc{test}},\mathcal{F}_{\textsc{test}})$ denote the graph during inference, where $\mathcal{G}_{\textsc{train}}\cup\mathcal{G}_{\textsc{test}}=\mathcal{G}_{\textsc{o}}$. The task falls into three regimes, based on the overlap between $\mathcal{G}_{\textsc{train}}$ and $\mathcal{G}_{\textsc{test}}$:
  • Transductive, in which the sets of entities and relation identifiers are fully shared ($\mathcal{G}_{\textsc{train}}=\mathcal{G}_{\textsc{test}}$). Traditional KG embedding models like TransE Bordes et al. (2013), RotatE Sun et al. (2019), and ComplEx Trouillon et al. (2016) learn embeddings for all entities and relations under this assumption. These models cannot predict links that involve entities or relations not observed and used for generating the embedding.
  • Partially Inductive. In this regime, all relation identifiers are known at training time ($\mathcal{R}_{\textsc{train}}=\mathcal{R}_{\textsc{test}}$), but not all the entities ($\mathcal{E}_{\textsc{train}}\subset\mathcal{E}_{\textsc{test}}$). Inductive Graph Neural Networks (GNNs) or rule mining methods like NBFNet Zhu et al. (2022) and GraIL Teru et al. (2020) learn to generalize in order to predict links involving entities from $\mathcal{E}_{\textsc{t}}\setminus\mathcal{E}_{\textsc{o}}$. However, they cannot predict links involving unseen relation identifiers.
  • Fully Inductive. In this most general regime, neither entities nor relations are fully known during training ($\mathcal{E}_{\textsc{train}}\subset\mathcal{E}_{\textsc{test}}$ and $\mathcal{R}_{\textsc{train}}\subset\mathcal{R}_{\textsc{test}}$). This regime best reflects the link prediction capabilities expected from a foundation model, namely, the ability to apply its pre-trained knowledge to entirely new, previously unseen KGs at test time. KGFMs, such as Ultra (Galkin et al., 2024), Trix (Zhang et al., 2025b) and Motif (Huang et al., 2025) learn widely transferable structural patterns.
@@
To the best of our knowledge, the state-of-the-art in link prediction has either considered textual attributions but ignored the fully inductive regime Yuan et al. (2025); Yao et al. (2019); Wei et al. (2023); Xu et al. (2021), or it has considered the fully inductive regime but has ignored text information Lee et al. (2023); Zhu et al. (2022); Teru et al. (2020). Our main hypothesis is that leveraging textual attributes can enhance fully inductive link prediction. The closest work to ours is ProLINK Wang et al. (2024), focused on working in the low-resource setting where the query relations occur sparsely in $\mathcal{G}_{\textsc{o}}$. ProLINK prompts LLMs to predict potential entity connections for unseen relations during testing, aiming to mitigate data sparsity. However, both their motivation and final evaluation setup differ from ours.
@@ 3 Foundation Models for KGs
Ultra. Ultra (Galkin et al., 2024) is a KGFM designed for fully inductive link prediction, capable of zero-shot generalization to unseen KGs. However, it does not exploit textual attributions. Ultra learns transferable structural patterns based on how relations interact within a graph, but it does not learn embeddings for specific relation types. Theoretically grounded by the principle of double equivariance Zhou et al. (2025), Ultra lifts the training and test graph $\mathcal{G}_{\textsc{o}}$ to a higher-level structure, which we refer to as ${\mathcal{G}_{R}^{\textsc{str}}}$. In ${\mathcal{G}_{R}^{\textsc{str}}}$, each node corresponds to a unique relation type or its inverse from $\mathcal{G}_{\textsc{o}}$. Edges in ${\mathcal{G}_{R}^{\textsc{str}}}$ capture fundamental structural interactions between pairs of relations in $\mathcal{G}_{\textsc{o}}$, categorized into head-to-head (h2h), tail-to-head (t2h), head-to-tail (h2t), tail-to-tail (t2t). ${\mathcal{G}_{R}^{\textsc{str}}}$ represents the transferable structural aspect, independent of the specific entity or relation vocabulary. Ultra utilizes a GNN, specifically NBFNet (Zhu et al., 2022), over ${\mathcal{G}_{R}^{\textsc{str}}}$, leveraging a labeling trick Zhang et al. (2021) to obtain conditional relation representations. These conditional relation representations are used as input features for a second NBFNet operating on the original graph $\mathcal{G}_{\textsc{o}}$ to perform the final link prediction. By learning relative structural patterns via ${\mathcal{G}_{R}^{\textsc{str}}}$ and conditioning representations on the query, Ultra avoids learning fixed relation embeddings, enabling its zero-shot transfer capabilities. Trix (Zhang et al., 2025b) and Motif (Huang et al., 2025) are follow-up works that enhanced Ultra’s representation power by leveraging extra memory and compute.
What do they lack? A key limitation of current KGFMs like Ultra, as discussed in Section 1, is their sole reliance on graph structure, often neglecting or underutilizing the textual attributions commonly associated with entities and relations in most real-world KGs. These textual attributions, especially the semantics encoded in relation names or descriptions, can be crucial for accurate and generalizable reasoning but are often overlooked by structure-focused methods (Cornell et al., 2022; Alam et al., 2024). The generalization capabilities of LLMs have been widely recognized and leveraged in other domains Li et al. (2024a), yet their potential to enrich KGFMs by interpreting textual attributes remains largely unexplored.
@@ 4 Semma: Integrating Structure and Textual Semantics
Our approach. To address the established desiderata, we propose the Semma architecture (Fig. 2), which explicitly separates structural processing (as in Ultra) from a module dedicated to processing relation’s textual semantics derived from LLMs in parallel. This dual-module approach provides modularity, allowing the text processing module to be deactivated when the textual attributions are noisy or unhelpful, thus maintaining robust performance ( 2 ). When suitable, the modules merge to effectively leverage structural and textual modalities ( 1 ).
We modularize our pipeline into three independent components: utilizing LLMs to generate semantically rich text for each relation (Section 4.1), embedding the extracted text attributions into a vector space (Section 4.2), and using the embeddings to incorporate the semantics of text attributions into the prediction pipeline (Section 4.3).
@@ 4.1 Extracting textual semantic information from LLMs
How do we extract this information from LLMs? We employ zero-shot prompting with illustrative examples and strict output constraints. This guides the LLM to produce consistent, reliable, and semantically meaningful text suitable for embedding across diverse relation inputs. More details about the prompt are provided in Appendix A.
@@ 4.2 Extracting embeddings from LLMs
Once we have the enriched textual representations from Section 4.1, we encode them into vector embeddings using transformers that capture their semantic meaning. These embeddings are the primary input for the text processing module.
Which text representation yields the most suitable embeddings? The process in Section 4.1 yields multiple textual candidates for each relation: (1) the original identifier (rel_name), (2) the LLM-generated cleaned name (llm_rel_name), and (3) the LLM-generated textual description (llm_rel_desc). Backed by the theory of concepts in the representation space (Park et al., 2025), we also create combined_sum, where the embeddings of all the above are combined through vector addition, and combined_avg, where they are averaged in the vector space. We denote the text embedding of a relation $r$ by $\mathbf{\tau}_{r}$, which can be obtained by one of the above-mentioned approaches. We evaluate Semma with all of the mentioned five variants, with more details in Section 5.3.
How do we embed inverse relations? For each relation $r$, Ultra generates its virtual inverse relation, denoted by $r^{-1}$, without its ground-truth text. Since we cannot embed the inverse relations without text (except in llm_rel_desc), we require a method to derive their semantic embeddings. Intuitively, if the original relation embedding signifies a transformation in a certain conceptual direction in the embedding space, its inverse should logically point the other way Sun et al. (2019). So, given a relation’s embedding $\mathbf{\tau}_{r}$, we generate the inverse relation’s embedding $\mathbf{\tau}_{r^{-1}}$ by rotating the original vector by 180 degrees to address this, i.e., $\mathbf{\tau}_{r^{-1}}=(\mathbf{I}-2\frac{\mathbf{\tau}_{r}\mathbf{\tau}_{r}^{T}}{\|\mathbf{\tau}_{r}\|^{2}})\mathbf{\tau}_{r}$ which simplifies to $\mathbf{\tau}_{r^{-1}}=-\mathbf{\tau}_{r}$. While theoretically imperfect for symmetric relations (which we observe to be rare in practice), this method is a reasonable heuristic and proves empirically effective (seen in Table 5).
@@ 4.3 Utilizing the embeddings
The final stage integrates the textual relation embeddings (from Section 4.2) with the structural information processed by Semma in parallel.
How do we represent inter-relation semantics? We construct a Textual Relation Graph (${\mathcal{G}_{R}^{\textsc{text}}}$) where nodes are relations and weighted edges (with weights $\omega$) represent the cosine similarity between their textual embeddings, to explicitly model semantic proximity between relations. An edge in ${\mathcal{G}_{R}^{\textsc{text}}}$ is denoted by a triple $(u,\omega,v)$ where $u,v$ are two relations with $\omega=\cos(\tau_{u},\tau_{v})$. For efficiency and to avoid oversmoothing of representations, we filter the edges ${\mathcal{G}_{R}^{\textsc{text}}}$ based on edge weights ($\omega$) using sampling methods like choosing the Top-xand thresholding Putra and Tokunaga (2017).
@@
$\mathbf{h}_{u}^{(t+1)}=\textsc{Upd}\left(\mathbf{h}_{u}^{(t)},\bigoplus\nolimits_{\begin{subarray}{c}v\in\mathcal{N}(u)\\ (u,r^{\prime},v)\in{\mathcal{G}_{R}^{\textsc{str}}}\end{subarray}}\textsc{Msg}(\mathbf{h}_{v}^{(t)},\mathbf{e}_{r^{\prime}})\right)\end{aligned}$
$\mathbf{z}_{u}^{(t+1)}=\textsc{Upd}\left(\mathbf{z}_{u}^{(t)},\bigoplus\nolimits_{\begin{subarray}{c}v\in\mathcal{N}(u)\\ (u,\omega,v)\in{\mathcal{G}_{R}^{\textsc{text}}}\end{subarray}}\textsc{Msg}(\mathbf{z}_{v}^{(t)},\mathbf{\omega})\right)\end{aligned}$
@@
where $\oplus$ is the concatenation operator and $\alpha$ is a hyperparameter that disables the text processing module if set to zero ($\alpha=1$ by default). This is then passed onto the Entity NBFNet for final prediction. We use the same training objective as Ultra for the whole pipeline.

### v2

@@ Abstract
Knowledge Graph Foundation Models (KGFMs) have shown promise in enabling zero-shot reasoning over unseen graphs by learning transferable patterns. However, most existing KGFMs rely solely on graph structure, overlooking the rich semantic signals encoded in textual attributes. We introduce Semma, a dual-module KGFM that systematically integrates transferable textual semantics alongside structure. Semma leverages Large Language Models (LLMs) to enrich relation identifiers, generating semantic embeddings that subsequently form a textual relation graph, which is fused with the structural component. Across 54 diverse KGs, Semma outperforms purely structural baselines like Ultra in fully inductive link prediction. Crucially, we show that in more challenging generalization settings, where the test-time relation vocabulary is entirely unseen, structural methods collapse while Semma is 22x more effective. Our findings demonstrate that textual semantics are critical for generalization in settings where structure alone fails, highlighting the need for foundation models that unify structural and linguistic signals in knowledge reasoning.
@@ 1 Introduction
Analogous to Large Language Models (LLMs) that learn complex correlations between tokens from a large corpus and apply them in previously unseen contexts Zhao et al. (2025), related work has started to investigate Knowledge Graph Foundation Models (KGFMs) to learn complex reasoning capabilities from training knowledge graphs and apply them to previously unseen test knowledge graphs Galkin et al. (2024a); Zhang et al. (2025b); Huang et al. (2025).
@@
Generalizing within knowledge graph embedding methods to transfer embeddings from one or multiple knowledge graphs with millions of nodes to an unseen knowledge graph with largely distinct nodes can be hindered by a low signal-to-noise ratio Zhao et al. (2021). However, we observe that knowledge graphs tend to have a huge number of nodes, but a much smaller number of highly informative relations. For example, Wikidata Vrandecic and Krötzsch (2014) contains over 110 million nodes but only 12,681 relations as of May 19, 2025. This disparity indicates that the compact vocabulary of relations, rather than the vast and often graph-specific set of entities, provides a more robust transferable signal for generalization.
@@
As illustrated in Figure˜1, our approach extends Ultra (Galkin et al., 2024a) by constructing: (i) a structural relation graph, which represents patterns that generalize from graph structures of the training graphs, similar to Ultra , and (ii) a textual relation graph, which represents patterns that are found by prompting an LLM to generate semantically rich descriptions for each relation identifier and/or attributed labels. Together, the two relation graphs allow for predicting high-quality links that would not be found by existing KGFM methods. We also make our codebase public.  [^1] https://github.com/arvindh75/semma In summary, our contributions are:
@@ 2 Preliminaries
Given the observed graph $\mathcal{G}_{\textsc{o}}$, link prediction aims to infer missing true triples $\mathcal{F}_{\textsc{t}}\setminus\mathcal{F}_{\textsc{o}}$ using triples with either the tail or head entity masked (called queries), denoted as $(h,r,?)$ and $(?,r,t)$. The latter can equivalently be written as $(t,r^{-1},?)$, where $r^{-1}$ denotes the inverse relation of $r$. For learning, we split the observed KG into a training graph $\mathcal{G}_{\textsc{train}}=(\mathcal{E}_{\textsc{train}},\mathcal{R}_{\textsc{train}},\mathcal{F}_{\textsc{train}})$ and a test graph $\mathcal{G}_{\textsc{test}}=(\mathcal{E}_{\textsc{test}},\mathcal{R}_{\textsc{test}},\mathcal{F}_{\textsc{test}})$, such that $\mathcal{G}_{\textsc{train}}\cup\mathcal{G}_{\textsc{test}}=\mathcal{G}_{\textsc{o}}$. The task falls into three regimes, based on the overlap between $\mathcal{G}_{\textsc{train}}$ and $\mathcal{G}_{\textsc{test}}$:
  • Transductive. In this regime, the training graph is identical to the test graph ($\mathcal{G}_{\textsc{train}}=\mathcal{G}_{\textsc{test}}$). Traditional KG embedding models like TransE Bordes et al. (2013), RotatE Sun et al. (2019), and ComplEx Trouillon et al. (2016) learn embeddings for all entities and relations under this assumption. These models cannot predict links that involve entities or relations not observed and used for generating the embedding.
  • Partially Inductive. In this regime, all relation identifiers are known at training time ($\mathcal{R}_{\textsc{train}}=\mathcal{R}_{\textsc{test}}$), but not all the entities ($\mathcal{E}_{\textsc{train}}\subset\mathcal{E}_{\textsc{test}}$). Inductive Graph Neural Networks (GNNs) or rule mining methods like NBFNet Zhu et al. (2022) and GraIL Teru et al. (2020) learn to generalize in order to predict links involving entities from $\mathcal{E}_{\textsc{test}}\setminus\mathcal{E}_{\textsc{train}}$. However, they cannot predict links involving unseen relation identifiers.
  • Fully Inductive. In this most general regime, neither entities nor relations are fully known during training ($\mathcal{E}_{\textsc{train}}\subset\mathcal{E}_{\textsc{test}}$ and $\mathcal{R}_{\textsc{train}}\subset\mathcal{R}_{\textsc{test}}$). This regime best reflects the link prediction capabilities expected from a foundation model, namely, the ability to apply its pre-trained knowledge to entirely new, previously unseen KGs at test time. KGFMs, such as Ultra (Galkin et al., 2024a), Trix (Zhang et al., 2025b) and Motif (Huang et al., 2025) learn widely transferable structural patterns.
@@
To the best of our knowledge, the state-of-the-art in link prediction has either considered textual attributions but ignored the fully inductive regime Yuan et al. (2025); Yao et al. (2019); Wei et al. (2023); Xu et al. (2021), or it has considered the fully inductive regime but has ignored text information Lee et al. (2023); Zhu et al. (2022); Teru et al. (2020). Prior works like BLP Daza et al. (2021), StatiK Markowitz et al. (2022), and SimKGC Wang et al. (2022) provide early empirical evidence that leveraging textual descriptions of entities and relations allows us to learn transferable representations for inductive prediction tasks without additional training. Building on this, our main hypothesis is that leveraging textual attributes can enhance fully inductive link prediction. The closest work to ours is ProLINK Wang et al. (2024), focused on working in the low-resource setting where the query relations occur sparsely in $\mathcal{G}_{\textsc{o}}$. ProLINK prompts LLMs to predict potential entity connections for unseen relations during testing, aiming to mitigate data sparsity. However, both their motivation and final evaluation setup differ from ours.
@@ 3 Foundation Models for KGs
Ultra. Ultra (Galkin et al., 2024a) is a KGFM designed for fully inductive link prediction, capable of zero-shot generalization to unseen KGs. However, it does not exploit textual attributions. Ultra learns transferable structural patterns based on how relations interact within a graph, but it does not learn embeddings for specific relation types. Theoretically grounded by the principle of double equivariance Zhou et al. (2025), Ultra lifts the training and test graph $\mathcal{G}_{\textsc{o}}$ to a higher-level structure, which we refer to as $\mathcal{G}_{R}^{\textsc{str}}$}. In $\mathcal{G}_{R}^{\textsc{str}}$}, each node corresponds to a unique relation type or its inverse from $\mathcal{G}_{\textsc{o}}$. Edges in $\mathcal{G}_{R}^{\textsc{str}}$} capture fundamental structural interactions between pairs of relations in $\mathcal{G}_{\textsc{o}}$, categorized into head-to-head (h2h), tail-to-head (t2h), head-to-tail (h2t), tail-to-tail (t2t). $\mathcal{G}_{R}^{\textsc{str}}$} represents the transferable structural aspect, independent of the specific entity or relation vocabulary. Ultra utilizes a GNN, specifically NBFNet (Zhu et al., 2022), over $\mathcal{G}_{R}^{\textsc{str}}$}, leveraging a labeling trick Zhang et al. (2021) to obtain conditional relation representations. These conditional relation representations are used as input features for a second NBFNet operating on the original graph $\mathcal{G}_{\textsc{o}}$ to perform the final link prediction. By learning relative structural patterns via $\mathcal{G}_{R}^{\textsc{str}}$} and conditioning representations on the query, Ultra avoids learning fixed relation embeddings, enabling its zero-shot transfer capabilities. Trix (Zhang et al., 2025b) and Motif (Huang et al., 2025) are follow-up works that enhanced Ultra’s representation power by leveraging extra memory and compute.
What do they lack? A key limitation of current KGFMs like Ultra, as discussed in Section˜1, is their sole reliance on graph structure, often neglecting or underutilizing the textual attributions commonly associated with entities and relations in most real-world KGs. These textual attributions, especially the semantics encoded in relation names or descriptions, can be crucial for accurate and generalizable reasoning but are often overlooked by structure-focused methods (Cornell et al., 2022; Alam et al., 2024). The generalization capabilities of LLMs have been widely recognized and leveraged in other domains Li et al. (2024a), yet their potential to enrich KGFMs by interpreting textual attributes remains largely unexplored.
@@ 4 Semma: Integrating Structure and Textual Semantics
Our approach. To address the established desiderata, we propose Semma (Figure˜2), which explicitly separates structural processing (as in Ultra) from a module dedicated to processing relations’ textual semantics derived from LLMs in parallel. This dual-module approach provides modularity, allowing the text processing module to be deactivated when the textual attributions are noisy or unhelpful, thus maintaining robust performance ( 2 ). When suitable, the modules merge to effectively leverage structural and textual modalities ( 1 ).
We modularize our pipeline into three independent components: utilizing LLMs to generate semantically rich text for each relation (Section˜4.1), embedding the extracted text attributions into a vector space (Section˜4.2), and using the embeddings to incorporate the semantics of text attributions into the prediction pipeline (Section˜4.3).
@@ 4.1 Extracting textual semantic information from LLMs
How do we extract this information from LLMs? We use a zero-shot prompt that provides the LLM with the relation’s identifier, example triples from the KG for context, and strict output constraints. This method ensures the generation of consistent and semantically meaningful text suitable for embedding. Note that we only assume LLM outputs are beneficial but not necessarily optimal in all contexts. More details about the prompt are provided in Appendix˜A.
@@ 4.2 Extracting embeddings from LLMs
Once we have the enriched textual representations from Section˜4.1, we encode them into vector embeddings using transformers that capture their semantic meaning. These embeddings are the primary input for the text processing module.
Which text representation yields the most suitable embeddings? The process in Section 4.1 yields multiple textual candidates for each relation: (1) the original identifier (rel_name), (2) the LLM-generated cleaned name (llm_rel_name), and (3) the LLM-generated textual description (llm_rel_desc). Backed by the theory of concepts in the representation space (Park et al., 2025), we also create combined_sum, where the embeddings of all the above are combined through vector addition, and combined_avg, where they are averaged in the vector space. We denote the text embedding of a relation $r$ by $\mathbf{\tau}_{r}$, which can be obtained by one of the above-mentioned approaches. We evaluate Semma with all of the mentioned five variants, with more details in Section˜5.3.
How do we embed inverse relations? For each relation $r$, Ultra generates its virtual inverse relation, denoted by $r^{-1}$, without its ground-truth text. Since we cannot embed the inverse relations without text (except in llm_rel_desc), we require a method to derive their semantic embeddings. Intuitively, if the original relation embedding signifies a transformation in a certain conceptual direction in the embedding space, its inverse should logically point the other way Sun et al. (2019). So, given a relation’s embedding $\mathbf{\tau}_{r}$, we generate the inverse relation’s embedding $\mathbf{\tau}_{r^{-1}}$ by rotating the original vector by 180 degrees to address this, i.e., $\mathbf{\tau}_{r^{-1}}=(\mathbf{I}-2\frac{\mathbf{\tau}_{r}\mathbf{\tau}_{r}^{T}}{\|\mathbf{\tau}_{r}\|^{2}})\mathbf{\tau}_{r}$ which simplifies to $\mathbf{\tau}_{r^{-1}}=-\mathbf{\tau}_{r}$. While theoretically imperfect for symmetric relations (which we observe to be rare in practice), this method is a reasonable heuristic and proves empirically effective (seen in Table˜5).
@@ 4.3 Utilizing the embeddings
The final stage integrates the textual relation embeddings (from Section˜4.2) with the structural information processed by Semma in parallel.
How do we represent inter-relation semantics? We construct a Textual Relation Graph ($\mathcal{G}_{R}^{\textsc{text}}$}) where nodes are relations and weighted edges (with weights $\omega$) represent the cosine similarity between their textual embeddings, to explicitly model semantic proximity between relations. An edge in $\mathcal{G}_{R}^{\textsc{text}}$} is denoted by a triple $(u,\omega,v)$ where $u,v$ are two relations with $\omega=\cos(\tau_{u},\tau_{v})$. For efficiency and to avoid oversmoothing of representations, we filter the edges $\mathcal{G}_{R}^{\textsc{text}}$} based on edge weights ($\omega$) using sampling methods like choosing the Top-x% and thresholding Putra and Tokunaga (2017).
@@
$\begin{aligned} \mathbf{h}_{u}^{(t+1)}=\textsc{Upd}\left(\mathbf{h}_{u}^{(t)},\bigoplus\nolimits_{\begin{subarray}{c}v\in\mathcal{N}(u)\\ (u,r^{\prime},v)\in{\mathcal{G}_{R}^{\textsc{str}}}\end{subarray}}\textsc{Msg}(\mathbf{h}_{v}^{(t)},\mathbf{e}_{r^{\prime}})\right)\end{aligned}$
$\begin{aligned} \mathbf{z}_{u}^{(t+1)}=\textsc{Upd}\left(\mathbf{z}_{u}^{(t)},\bigoplus\nolimits_{\begin{subarray}{c}v\in\mathcal{N}(u)\\ (u,\omega,v)\in{\mathcal{G}_{R}^{\textsc{text}}}\end{subarray}}\textsc{Msg}(\mathbf{z}_{v}^{(t)},\mathbf{\omega})\right)\end{aligned}$
@@
where $\oplus$ is the concatenation operator and $\alpha$ is a hyperparameter that disables the text processing module if set to zero ($\alpha=1$ by default). This is then passed on to the Entity NBFNet for final prediction. We use the same training objective as Ultra for the whole pipeline. The primary overhead introduced by Semma is a one-time LLM enrichment step, which has a computational complexity of $O(|\mathcal{R}_{\textsc{test}}|)$, scaling linearly with the number of relations in $\mathcal{G}_{\textsc{test}}$. More details in Appendix˜D.